What is the range of the function f(x) = -x - 3, when the domain is {-1, 0, 4, 7}?
Answer by rapaljer(4551)

Substitute the values of x to see what you get for f(x):
f(x) = -x - 3
f(-1) = -(-1) - 3 = -2
f(0) = -0 - 3= -3
f(4) = -4 - 3 =-7
f(7) = -7-3 = -10

Range is the set of all y-values, or in this case, all f(x) values, so
Range = { -2, -3, -7, -10}

R^2 at SCC